OpenStudy (anonymous):

whats the value of this equation.? 4x-2=3

OpenStudy (anonymous):

4x= 3+2 4x= 5 x= 5/4 x=1.25.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

x=5 over 4

OpenStudy (anonymous):

yup, just follow nikko's answer ^^

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

niko is true...